Moe earns 11% more than her husband Joe. Together they make $ 62000 per year. What is Joe’s annual salary?
I tried finding 11% of 62000 which results in 6820 and then subtracting from the total salary. Then from the subtracted total i tried halving the total to find Joe's salary.

Answer by Maths68(1474)

Let
Salary of Joe = j
Salary of Moe = m
Given
Sum of both salaries (Joe + Moe) = 62000
11% of j = j* 11% = j * 11/100 = 11j/100
Moe earns 11% more than her husband Joe
Moe Salary = Joe's salary + 11% of Joe's salary
Moe Salary = j + 11j/100
Joe salary + Moe salary = Sum of salaries
j + j + 11j/100 = 62000
j + j + 11j/100 = 62000
2j + 11j/100 = 62000
Multiply by 100 both sides of above equation
100*2j + 100*(11j)/100] = 100*62000
200j+11j=6200000
211j=6200000
Divide by 211 both sides of above equation
211j/211=6200000/211
j=29383.8862559
Moe Salary = j + 11j/100
Put the value of j in above equation
Moe Salary = 29383.8862559+11(29383.8862559)/100
Moe Salary = 29383.8862559+323222.748814/100
Moe Salary = 29383.8862559+3232.22748814
Moe Salary = 32616.11374

Joe salary + Moe salary = Sum of salaries
29383.8862559+32616.11374 = 62000
62000=62000